# The Quantum Measurement Problem: A Review of Recent Trends

### v1

@@ 5 Facing the Problem of Outcomes
In the previous section, we have seen that decoherence clarifies a significant portion of the Measurement Problem. However, the problem of outcomes is still open: After the basis is chosen and quantum superpositions are suppressed, the system still remains on a mixture of possible outcomes. Decoherence does not tell how and why only one of these outcomes is measured. At this point, standard quantum mechanics interpretation claims the occurrence of an additional dynamic step, the collapse:
We can illustrate this point with an example of a photoexcited molecule undergoing internal conversion Figure 4. Light initially promotes the molecule to an excited electronic state. The nuclear wavepacket evolves on the potential energy surface of the excited state until it reaches a degeneracy region with the ground state potential energy surface. There, quantum coherence builds between the two states, creating a state superposition. The dephasing between the wavepacket components in each state counteracts, causing decoherence. All events up to this point, including decoherence, can be simulated by propagating the Schrödinger equation for the molecule [34]. If the molecule is measured, it will either be in the ground or excited state. However, our simulation can only tell the probabilities of these outputs.
Standard quantum mechanics assumes a wave function collapse occurred. Nevertheless, the existence of collapse is far from consensual. It is regarded as a superfluous hypothesis [35] in several quantum mechanical interpretations.
In the following subsections, we will quickly dive into a few classes of potential solutions that represent the current debate on the Measurement Problem, with little attention to philosophical and historical aspects. More comprehensive approaches can be found in Refs.[6], [32, Introduction & Ch 11], [36, 37].
Before that, we should clarify the concepts we are dealing with. In literature, it is common to use decoherence and collapse concepts interchangeably. We should avoid it, as these phenomena may, in principle, be completely independent and occur in different time scales.
As stated in Eq. (21), decoherence acts on the off-diagonal terms of the system’s reduced density matrix. Collapse, however, zeroes the diagonal terms except for one element. We can be a bit more flexible with this definition. In principle, a quantum theory including collapse could go as far as to provide a diagonal density matrix. This means we would be left in a situation of statistical uncertainty: the collapse occurred, but the theory only informed us of the probability of each outcome. We could call such a case an epistemic collapse. If, however, the theory zeroes the diagonal (except one) elements, then we would have an ontic collapse.
@@ 5.1 Many-Worlds Interpretation
The lack of information on the outcomes is not a problem for some interpretations. For instance, the Many-Worlds Interpretation (or Relative State Interpretation or, still, Everettian Quantum Mechanics) proposes that the remaining outcomes after decoherence correspond to independent universe branches [38]. Observers existing in each branch perceive different outcomes because their quantum state, entangled with the molecular quantum state, also branches. In one branch, the combined molecule-observer quantum state may be and in another, . There are no branches with superpositions of both states because of decoherence toward the environment surrounding the system and the observer suppressing them [39]. Thus, the Many-Worlds Interpretation posits the existence of a single quantum universe composed of non-communicating classical branches.
For the Many-Worlds Interpretation, the quantum state is a complete description of reality (no hidden variables are required), exclusively evolving according to the Schrödinger Equation (collapse is only an illusion caused by the universe branching) [40, 41]. Thus, the Many-Worlds Interpretation is consistent with standard quantum mechanics and requires no theory modification. However, it remains an interpretation rather than a theory until it gains experimental testability. A significant challenge is the technical infeasibility of maintaining and measuring a quantum system in a coherent superposition of orthogonal states, especially for large systems. This difficulty explains why superpositions of macroscopic states, such as and , cannot be observed directly due to rapid decoherence. Susskind and Aaronson demonstrated this limitation using methods from quantum computational complexity, naming it the quantum-hard problem [42, 43].
@@
The Many-Worlds Interpretation may seem exotic to non-specialists due to its exuberant ontology. Despite this perception, it is a serious contender among the proposed solutions to the Measurement Problem. It is a popular interpretation among physicists, with prominent supporters like Schlosshauer [20], Carroll [40, 41, 44], Zurek [39], Zeh [35], and Deutch promoting different variants. However, it also faces outstanding critics, like Penrose [47] and Albert. For Landsmann, the Many-Worlds Interpretation would be “acceptable only if truly everything else has failed” [32, Ch 11].
@@ 5.2.1 Relational Interpretation
Rovelli draws attention to the fact that this state dependence on a specific observer is not new in physics. It is analogous, for instance, to a velocity dependence on the particular reference frame where it’s measured. The novelty in quantum mechanics is that the non-commutative nature of quantum variables implies that not all properties can be sharply defined simultaneously [50]. This limits the completeness of information derivable from a single interaction.
@@
Giacomini, Castro-Ruiz, and Brukner have recently extended relational ideas beyond Rovelli’s Relational Quantum Mechanics by incorporating Quantum Reference Frames, QRF [54]. In their approach, a QRF is simply a quantum system elevated to the role of a reference frame so that its degrees of freedom define the origin relative to which one describes other systems. By extending canonical transformations to the quantum domain, Giacomini et al. show how usual frame changes (translations, for example) remain valid even if the reference frame itself is in a superposition or entangled state. This leads to the relativity of notions like superposition and entanglement. If a given particle is in a spatial superposition from the lab’s viewpoint, then from the particle’s viewpoint, the lab may itself appear in a corresponding superposition.
This perspective may offer new insights into the Measurement problem, particularly in Wigner’s friend scenarios, where the possibility of conflicting accounts among observers becomes natural if each observer constitutes its own QFR [54, 55]. In Section 5.3.2, we discuss how Giacomini and Brukner use QFRs to challenge Pennrose’s conjecture on gravity-induced collapse [56].
@@ 5.2.2 Information-Theoretic and QBism Views
Another epistemic perspective comes from Quantum Information Theory, which treats quantum mechanics as an axiomatic framework akin to thermodynamics. Brukner and Zeilinger [51] argue that quantum mechanics should be understood as a theory of information rather than a theory of reality, where the fundamental postulates describe how information is processed rather than how physical systems evolve.
Recent experimental results reinforce this perspective. Spegel-Lexne and co-authors experimentally demonstrated the equivalence of entropic uncertainty and wave-particle duality [57], reinforcing that quantum measurement can be understood through informational constraints rather than wave function collapse.
The epistemic perspective in Quantum Information Theory aligns with Quantum Bayesianism (QBism), which takes a radically subjective approach: the quantum state does not represent an objective reality but rather an observer’s personal beliefs about measurement outcomes [52]. When a measurement is made—such as determining if the molecule is in the state—the observer updates the quantum state to reflect the new information [52, 58].
However, Information-Theoretic interpretations, particularly QBism, face several challenges. The Pusey-Barret-Rudolph (PBR) theorem [59] states that epistemic interpretations of the quantum states cannot entirely reproduce the results of quantum mechanics. It has also been claimed that epistemic interpretations (at least in the QBism form) contradict Ozawa’s Intersubjectivity theorem, which poses that two observers jointly measuring the same system must obtain the same outcome [60, 61]. However, no-go theorems can always be criticized for their underlying hypotheses [62] (We come back to this point in Section 5.4.).
@@ 5.2.3 Indivisible Stochastic Process Theory
Barandes has recently proposed the Indivisible Stochastic Processes theory, where classical configurations of particles, fields, or qubits evolve via non-Markovian dynamics [53]. In this type of dynamics, which recovers all results from conventional quantum mechanics, transition probabilities cannot be estimated at intermediate times. In Barandes’ approach, wave functions are considered predictive tools only, and the Hilbert space of quantum mechanics is demoted to a convenient instrumentalist framework without an ontological basis.
Barandes draws a historical analogy, comparing the transition from instrumentalist frameworks in quantum mechanics to the Copernican revolution, where the instrumental Ptolemaic system of epicycles was replaced by a more fundamental description based on forces. Similarly, he advocates for a clear physical picture to supersede the abstract formalism of Hilbert space mechanics. Like the Relational Interpretation (Section 5.2.1), the Indivisible Stochastic Process formulation features a sparse ontology, emphasizing that properties are not well-defined between interactions but are instead contextualized through the system’s stochastic dynamics. By embedding quantum systems within generalized stochastic systems, this approach aligns with a realist philosophy, reinterpreting quantum systems as governed by dynamic interactions without requiring measurement to be treated as a distinct or special process.
@@ 5.3.1 Agnostic Objective Collapse
Diametraly opposed to Many-Worlds and Epistemic Interpretations, Objective Collapse Theories consider that collapse is a physical process, and that standard quantum mechanics is still incomplete [4, 5]. Thus, these theories attempt to change the Schrödinger Equation to account for the collapse. This subsection deals with agnostic objective collapse, that is, theories that don’t attribute a cause to the collapse. In the following subsection, we will look at gravity-induced collapse.
Most objective collapse theories predict that a system in quantum superposition randomly localizes in one of the outcomes given enough time [5, 63, 64, 65]. The quantum state is driven by a Stochastic Schrödinger Equation, such as the Gisin Equation [66, 67]
@@
Different objective collapse theories have been proposed, corresponding to specific choices of operators , including the Hamiltonian operator [68], position operators (GRW after the initials of the authors [69]) density number operators (CSL for Continuous Spontaneous Localization [64]), and mass density operators [70] (which is discussed in Section 5.3.2). These specific models are thoroughly reviewed in Ref. [4].
@@ 5.3.2 Gravity-Induced Objective Collapse
and is the gravitational constant. This collapse time can also be derived by computing the difference between Newtonian free fall accelerations in the space surrounding each geometry [72]. A direct application of Eq. (23) predicts it would take billions of years to collapse the wave function of an isolated molecule, but only seconds to observe collapse in a kg body (Figure 6) [1].
@@
The Diósi-Penrose model has attracted criticism on multiple fronts. Gao [73] raised several issues on the Penrose conjecture. His main point is that Penrose’s argument by analogy with the conventional quantum mechanical time-uncertainty principle is not rigorous enough. Moreover, he is also right to point out that although we commonly refer to Diósi-Penrose model, Diósi and Penrose theories are fundamentally distinct, only sharing, maybe by simple coincidence, the same collapse time. Giacomini and Bruckner [56] also challenged Penrose’s conjecture that quantum superposition conflicts with Einstein’s equivalence principle. Using the Quantum Reference Frames (QRF) formalism (see also 5.2.1), they showed that it’s possible to generalize the principle (or, more precisely, always to make the metric field locally Minkowskian) even in the presence of quantum superpositions of massive bodies.
Additionally, we performed atomistic simulations to show that the Diósi-Penrose collapse time reaches a minimum value independent of superposition displacement and, in some instances, can be surprisingly long [1]. Figurato et al. [74] has also independently concluded that the model could have too long collapse times for macroscopic bodies. Beyond these theoretical concerns, experimental tests have also challenged the model. Donadi et al. [75] and Arnquist et al. [76] conducted an underground experiment to test gravity-related collapse mechanisms. Their results placed stringent constraints on the model’s parameters.
Although all these points must be considered when assessing the validity of the Diósi-Penrose model, none of them directly falsify the model. And, as Thorne said, ”Penrose has a way of always being right, even when he does things in a strange way” [77].
Oppenheim has recently proposed another hypothesis that attributes wave function collapse to the influence of gravity. Indeed, his Postquantum Theory of Classical Gravity does not focus on the measurement problem and has a more ambitious goal of reconciling gravity and quantum mechanics [78]. Oppenheim’s central hypothesis is that gravity is a classical field. To ensure classical gravity can interact coherently with quantum matter fields, Oppenheim developed a Classical Quantum (CQ) Dynamics framework, which circumvents the usual obstacles encountered by semiclassical models, such as violation of the superposition principle due to nonlinearities in the Schrödinger equation [78, 79, 80].
The CQ dynamics of the hybrid state (where are the classical phase space coordinates) is expressed in terms of the general master equation [78]
@@
Like in modeling quantum open systems [82], Eq.(5.3.2) is not directly solved for but rather through an ensemble of stochastic trajectories (unraveling procedure) [83, 84]. In each trajectory, the quantum degrees of freedom undergo stochastic projection onto the eigenstates of the corresponding Lindblad operators, collapsing to a definite outcome in a manner that statistically recovers the Born rule. These quantum jumps induce corresponding phase space jumps in the classical degrees of freedom, representing a backreaction of the quantum system on the classical one. In addition to collapse, this equation also describes decoherence, leading to the suppression of quantum coherence over time. Figure 7 illustrates the quantum and classical stochastic jumps in a system composed of a qubit and a classical particle discussed in Ref.[83]
@@
Even though stochastic trajectories are introduced as a means to propagate the density evolution (5.3.2), Oppenheim and coauthors assign them ontological significance [83]. A stochastic trajectory represents the system collapsing to a specific state, just like it is also done in the other objective collapse theories that we discussed above. Oppenheim’s theory implies that collapse does not require measurement, as it occurs naturally due to interactions with a classical field [83].
@@
Gravity-induced wave function collapse may have profound implications for physics. If any of these theories are confirmed, it would not only close the Measurement Problem but also imply that gravity plays a unique role in nature, different from the other fundamental interactions. Moreover, it would also impact future efforts to reconcile general gravity and quantum mechanics, moving toward what Penrose calls gravitization of quantum mechanics [85].

### v2

@@ 5 Facing the Problem of Outcomes
In the previous section, we have seen that decoherence clarifies a significant portion of the Measurement Problem. However, the problem of outcomes is still open: After the basis is chosen and quantum superpositions are suppressed, the system still remains in a mixture of possible outcomes. Decoherence does not tell how and why only one of these outcomes is measured. At this point, standard quantum mechanics interpretation claims the occurrence of an additional dynamic step, the collapse:
We illustrate these different processes with an example of a photoexcited molecule undergoing internal conversion in Figure 4. Light initially promotes the molecule to an excited electronic state. The nuclear wavepacket evolves on the potential energy surface of the excited state until it reaches a degeneracy region with the ground state potential energy surface. There, quantum coherence builds between the two states, creating a state superposition. The dephasing between the wavepacket components in each state counteracts, causing decoherence. All events up to this point, including decoherence, can be simulated by propagating the Schrödinger equation for the molecule [41]. If the molecule is measured, it will either be in the ground or excited state. However, our simulation can only tell the probabilities of these outputs.
Standard quantum mechanics assumes that a wave function collapse occurred. Nevertheless, the existence of collapse is far from consensual. It is regarded as a superfluous hypothesis [42] in several quantum mechanical interpretations.
In the following subsections, we will dive into a few classes of potential solutions that represent the current debate on the Measurement Problem, with little attention to philosophical and historical aspects. More comprehensive approaches can be found in Refs.[4], [39, Introduction & Ch 11], [43, 44].
Before that, we should clarify the concepts we are dealing with. In literature, decoherence and collapse concepts are sometimes used interchangeably. We should avoid it, as these phenomena may, in principle, be completely independent and occur on different time scales.
As stated in Eq. (25), decoherence acts on the off-diagonal terms of the system’s reduced density matrix. Collapse, however, zeroes the diagonal terms except for one element of the entire density matrix (including the observer), which is not represented in Eq. (25). We can be a bit more flexible with this definition. In principle, a quantum theory including collapse could go as far as to provide a diagonal density matrix. This means we would be left in a situation of statistical uncertainty: the collapse occurred, but the theory only informed us of the probability of each outcome. We could call such a case an epistemic collapse. If, however, the theory zeroes the diagonal (except one) elements, then we would have an ontic collapse.
@@ 5.1 Many-Worlds Interpretation
The lack of information on the outcomes is not a problem for some interpretations. For instance, the Many-Worlds Interpretation (or Relative State Interpretation or, still, Everettian Quantum Mechanics) proposes that the remaining outcomes after decoherence correspond to independent universe branches [45]. Observers existing in each branch perceive different outcomes because their quantum state, entangled with the molecular quantum state, also branches. In one branch, the combined molecule-observer quantum state may be and in another, . No branches with superpositions of both states exist because of decoherence toward the environment (surrounding the system and the observer) suppressing them [46]. Thus, the Many-Worlds Interpretation posits the existence of a single quantum universe composed of non-communicating classical branches.
For the Many-Worlds Interpretation, the quantum state is a complete description of reality (no hidden variables are required), exclusively evolving according to the Schrödinger Equation (collapse is only an illusion caused by the universe branching) [47, 48]. Thus, the Many-Worlds Interpretation is consistent with standard quantum mechanics and requires no theory modification. However, it remains an interpretation rather than a theory until it gains experimental testability. A significant challenge is the technical infeasibility of maintaining and measuring a quantum system in a coherent superposition of orthogonal states, especially for large systems. This difficulty explains why superpositions of macroscopic states cannot be observed directly due to rapid decoherence. Susskind and Aaronson demonstrated this limitation using methods from quantum computational complexity, naming it the quantum-hard problem [49, 50].
@@
The Many-Worlds Interpretation may seem exotic to non-specialists due to its exuberant ontology. Despite this perception, it is a serious contender among the proposed solutions to the Measurement Problem. It is a popular interpretation among physicists, with prominent supporters like Schlosshauer [24], Carroll [47, 48, 51], Zurek [46], Zeh [42], and Deutsch promoting different variants. However, it also faces outstanding critics, like Penrose [54] and Albert. For Landsmann, the Many-Worlds Interpretation would be “acceptable only if truly everything else has failed” [39, Ch 11].
@@ 5.2.1 Relational Interpretation
Rovelli draws attention to the fact that this state dependence on a specific observer is not new in physics. It is analogous, for instance, to a velocity dependence on the particular reference frame where it is measured. The novelty in quantum mechanics is that the non-commutative nature of quantum variables implies that not all properties can be sharply defined simultaneously [57], limiting the completeness of information derivable from a single interaction.
@@
Giacomini, Castro-Ruiz, and Brukner have recently extended relational ideas beyond Rovelli’s Relational Quantum Mechanics by incorporating Quantum Reference Frames, QRF [61]. In their approach, a QRF is simply a quantum system elevated to the role of a reference frame so that its degrees of freedom define the origin relative to which one describes other systems. By extending canonical transformations to the quantum domain, Giacomini et al. show how usual frame changes (translations, for example) remain valid even if the reference frame itself is in a superposition or entangled state. Their approach leads to the relativity of notions like superposition and entanglement. If a given particle is in a spatial superposition from the lab’s viewpoint, then from the particle’s viewpoint, the lab may itself appear in a corresponding superposition.
This perspective may offer new insights into the Measurement problem, particularly in Wigner’s friend scenarios, where the possibility of conflicting accounts among observers becomes natural if each observer constitutes its own QFR [61, 62]. Section 5.3.2 discusses how Giacomini and Brukner use QFRs to challenge Pennrose’s conjecture on gravity-induced collapse [63].
@@ 5.2.2 Information-Theoretic and QBism Views
Another epistemic perspective comes from Quantum Information Theory, which treats quantum mechanics as an axiomatic framework akin to thermodynamics. Brukner and Zeilinger [58] argue that quantum theory is fundamentally about information, not underlying physical reality. Its postulates describe how information is processed, not how systems evolve in space and time.
Recent experimental results reinforce this perspective. Spegel-Lexne and co-authors demonstrated the equivalence of entropic uncertainty and wave-particle duality [64], suggesting that quantum measurement can be understood through informational constraints rather than wave function collapse.
A more radical variant of this informational approach is Quantum Bayesianism, or QBism [59]. QBism interprets the quantum state not as a system property but as a tool that reflects an agent’s personal beliefs (or more precisely, doxastic states) about the outcomes of their measurements. For instance, assigning the state to a molecule expresses an agent’s expectation about what they will experience upon interacting with it, not a claim about an objective feature of the molecule. When a measurement is made, the quantum state is updated to reflect the agent’s new information, akin to a Bayesian update of probabilities. From this perspective, collapse is not a physical transformation but the agent updating their beliefs in response to their own experience, which is the measurement outcome [59, 65].
Some no-go theorems are sometimes cited as challenging epistemic views, particularly the Pusey-Barrett-Rudolph (PBR) theorem [66]. The theorem shows that, under certain assumptions, quantum states cannot be interpreted merely as statistical knowledge about underlying physical states—that is, hidden variables (see Section 5.4). Instead, the quantum state must be part of the physical reality. However, this conclusion applies only to models that posit such hidden variables. Interpretations like QBism, which deny the existence of underlying ontic states altogether, fall outside the scope of the theorem [67, 68]. Rovelli has made similar clarifications for relational interpretations as well [69].
Another supposed challenge to QBism comes from Ozawa’s intersubjectivity theorem [70, 71], which claims that two observers jointly measuring a system must agree on its outcome. This has been interpreted as incompatible with QBism’s agent-centric framework. However, Schack has recently argued that this conclusion is mistaken and that QBism maintains consistency across observers without requiring objective outcomes in the traditional sense [72].
A common misunderstanding is to conflate QBism with older subjectivist interpretations, like that of London and Bauer [73], who suggested that collapse results from an observer becoming aware of an outcome. However, QBism takes a different stance: the measurement outcome is not something the agent discovers; it is the agent’s experience. No underlying real wave function remains untouched—it is a theory about actions and consequences, not objective ontologies [74].
By rejecting objective quantum states, QBism sidesteps many interpretative paradoxes, offering a consistent but radically subjective view in which quantum theory serves as a decision-making tool for agents.
@@ 5.2.3 Indivisible Stochastic Process Theory
Barandes has recently proposed the Indivisible Stochastic Processes theory, where classical configurations of particles, fields, or qubits evolve via non-Markovian dynamics [60, 75]. In this type of dynamics, which recovers all results from conventional quantum mechanics, transition probabilities cannot be estimated at intermediate times. In Barandes’ approach, wave functions are considered predictive tools only, and the Hilbert space of quantum mechanics is demoted to a convenient instrumentalist framework without an ontological basis.
Barandes draws a historical analogy, comparing the transition from instrumentalist frameworks in quantum mechanics to the Copernican revolution, where the instrumental Ptolemaic system of epicycles was replaced by a more fundamental description based on forces. Similarly, he advocates for a clear physical picture to supersede the abstract formalism of Hilbert space mechanics.
Unlike the Relational Interpretation (Section 5.2.1), the Indivisible Stochastic Process formulation posits a non-perspectival, objective reality: each system has a definite configuration that evolves stochastically in time. The ontology is well-defined at all times: each system has a definite configuration, even between interactions. What is sparse are the dynamical laws, since conditional probabilities connecting two moments in time are generally defined only when the earlier time is either the initial time or a special kind of interaction event (called a division event). This framework does not assume divisibility or Markovianity.
By embedding quantum systems within generalized stochastic systems, the Indivisible Stochastic Process formulation aligns with a realist philosophy, reinterpreting quantum systems as governed by dynamic interactions without requiring measurement to be treated as a distinct or special process.
@@ 5.3.1 Agnostic Objective Collapse
Opposed to Many-Worlds and Epistemic Interpretations, Objective Collapse Theories consider that collapse is a physical process, and that standard quantum mechanics is still incomplete [5, 8]. Thus, these theories attempt to change the Schrödinger Equation to account for the collapse. This subsection deals with agnostic objective collapse, that is, theories that don’t attribute a cause to the collapse. In the following subsection, we will look at gravity-induced collapse.
Most objective collapse theories predict that a system in quantum superposition randomly localizes in one of the outcomes given enough time [8, 76, 77, 78]. The quantum state is driven by a Stochastic Schrödinger Equation, such as the quantum-state diffusion equation [79, 80]
@@
Different objective collapse theories have been proposed, corresponding to specific choices of operators , including the Hamiltonian operator [82], position operators (GRW after the initials of the authors [83]) density number operators (CSL for Continuous Spontaneous Localization [77]), and mass density operators [84] (which is discussed in Section 5.3.2). These specific models are thoroughly reviewed in Ref. [5, 9].
@@ 5.3.2 Gravity-Induced Objective Collapse
and is the gravitational constant. This collapse time can also be derived by computing the difference between Newtonian free-fall accelerations in the space surrounding each geometry [86]. A direct application of Eq. (27) predicts it would take billions of years to collapse the wave function of an isolated molecule, but only seconds to observe collapse in a kg body (Figure 6) [1].
@@
The Diósi-Penrose model has attracted criticism on multiple fronts. Gao [87] raised several issues on the Penrose conjecture. His main point is that Penrose’s argument by analogy with the conventional quantum mechanical time-uncertainty principle is not rigorous enough. Moreover, he is also right to point out that although we commonly refer to the Di’osi-Penrose model, Diósi and Penrose theories are fundamentally distinct, arriving at the same collapse time through independent, though conceptually related, arguments. Diósi himself [88] notes their similarities and differences and suggests that further work is needed to clarify whether the two frameworks are compatible.
Giacomini and Bruckner [63] also challenged Penrose’s conjecture that quantum superposition conflicts with Einstein’s equivalence principle. Using the Quantum Reference Frames (QRF) formalism (see also 5.2.1), they showed that it’s possible to generalize the principle (or, more precisely, always to make the metric field locally Minkowskian) even in the presence of quantum superpositions of massive bodies.
Additionally, we performed atomistic simulations to show that the Diósi-Penrose collapse time reaches a minimum value independent of superposition displacement and, in some instances, can be surprisingly long [1]. Figurato et al. [89] has also independently concluded that the model could have too long collapse times for macroscopic bodies. Beyond these theoretical concerns, experimental tests have also challenged the model. Donadi et al. [90] and Arnquist et al. [91] conducted underground experiments to test gravity-related collapse mechanisms. Their results placed stringent constraints on the model’s parameters (see Section 5.3.3).
Although all these points must be considered when assessing the validity of the Diósi-Penrose model, none of them directly falsifies the model. And, as Thorne said, ”Penrose has a way of always being right, even when he does things in a strange way” [92].
Oppenheim has recently proposed another hypothesis that attributes wave function collapse to the influence of gravity. Indeed, his Postquantum Theory of Classical Gravity does not focus on the measurement problem and has a more ambitious goal of reconciling gravity and quantum mechanics [93]. Oppenheim’s central hypothesis is that gravity is a classical field. To ensure classical gravity can interact coherently with quantum matter fields, Oppenheim developed a Classical Quantum (CQ) Dynamics framework, which circumvents the usual obstacles encountered by semiclassical models, such as violation of the superposition principle due to nonlinearities in the equations of motion [93, 94, 95].
The CQ dynamics of the hybrid state (where represents the classical phase space coordinates) is expressed in terms of the general master equation [93]
@@
Like in modeling open quantum systems [96], Eq.(5.3.2) is not directly solved for but rather through an ensemble of stochastic trajectories (unraveling procedure) [97, 98]. In each trajectory, the quantum degrees of freedom undergo stochastic projection onto the eigenstates of the corresponding Lindblad operators, collapsing to a definite outcome in a manner that statistically recovers the Born rule. These quantum jumps induce corresponding phase space jumps in the classical degrees of freedom, representing a backreaction of the quantum system on the classical one. In addition to collapse, this equation also describes decoherence, leading to the suppression of quantum coherence over time. Figure 7 illustrates the quantum and classical stochastic jumps in a system composed of a qubit and a classical particle discussed in Ref.[97]
@@
Even though stochastic trajectories are introduced as a means to propagate the density evolution (5.3.2), Oppenheim and co-authors assign them ontological significance [97]. A stochastic trajectory represents the system collapsing to a specific state, just as it is also done in the other objective collapse theories discussed above. Oppenheim’s theory implies that collapse does not require measurement, as it occurs naturally due to interactions with a classical field [97].
@@
Gravity-induced wave function collapse may have profound implications for physics. If any of these theories are confirmed, it would not only close the Measurement Problem but also imply that gravity plays a unique role in nature, different from the other fundamental interactions. Moreover, it would also impact future efforts to reconcile general gravity and quantum mechanics, moving toward what Penrose calls the gravitization of quantum mechanics [99].
5.3.3 Experimental Assessment of Objective Collapse Theories
Experimental efforts to test objective collapse models, especially CSL, have gained remarkable traction in recent years [6, 9, 91, 100, 101]. One of the most direct approaches involves matter-wave interferometry with large molecules or nanoparticles, designed to probe the persistence of quantum coherence at mesoscopic scales [1]. Gasbarri et al. argue that the microgravity environment of space, with its long free-fall times and low noise, offers a unique setting for such experiments, enabling both interferometric and non-interferometric tests with unprecedented sensitivity to collapse effects [101].
Beyond interferometry, several non-interferometric techniques have been proposed to test collapse models without requiring explicit superpositions. Carlesso et al. reviewed platforms such as optomechanical resonators, ultra-cold atoms, and solid-state systems, where collapse-induced noise may manifest as excess heating or anomalous diffusion [6]. Other proposals investigate spontaneous X-ray emission from the random motion of charged particles, a distinctive prediction of CSL. These diverse strategies have steadily improved the empirical bounds on collapse parameters, ruling out wide ranges of the originally proposed parameter space for both CSL and Diósi-Penrose models.
Recent studies have expanded the scope of experimental constraints. Altamura et al. analyzed data from the LISA Pathfinder mission and found that rotational degrees of freedom could provide tighter bounds than translational motion in some scenarios [100]. In a related direction, Howl, Penrose, and Fuentes proposed an experiment using a Bose–Einstein condensate prepared in a spatial superposition to test gravitationally induced collapse [86].
As discussed in Section 5.3.2, the Diósi-Penrose model ties collapse to gravitational self-energy. Figurato et al. investigated whether the model predicts rapid enough collapse for macroscopic systems and showed that it does not always do so across the entire parameter range [89]. Meanwhile, an underground experiment at the Gran Sasso Laboratory searched for spontaneous radiation arising from collapse-induced diffusion, placing strong limits on the spatial resolution of mass density—enough to exclude the original parameter-free version of the Diósi-Penrose proposal [90].
These efforts signal a shift in the status of objective collapse theories from speculative constructs to empirically testable frameworks. Among them, CSL stands out for its tunable parameters and broad compatibility with experimental platforms, while the Diòsi-Penrose and Oppenheim models offer more rigid, gravity-motivated alternatives. For example, Angeli et al. recently demonstrated that if gravity is both classical and local, as in Oppenheim’s proposal (see Section 5.3.2), then it must induce diffusion in quantum systems, a phenomenon potentially observable with large-scale mechanical probes such as torsion pendulums [102].
Altogether, these developments highlight the growing entanglement of collapse models with experimental physics. Whether or not they ultimately capture physical reality, they provide concrete benchmarks for testing the limits of quantum theory and may yet point toward deeper structures beyond it.